# A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness

### v1

@@ 3 A Formal Falsification Framework for Theories of Consciousness
Such universal substitutions provide a means of a priori falsification. Essentially, substituting them into an experimental setup for the original would radically alter or complicate the predictions of a theory of consciousness, while the outward and immediate inferences (such as report or output) would not change at all.
@@
However, the Kleiner-Hoel dilemma has two horns. If the predictions of a theory of consciousness are strictly dependent on the experimental inferences observed during testing a theory of consciousness, then that theory is unfalsifiable.
Here and throughout, such strictly-dependent theories are referred to as “trivial” to clearly distinguish that the source of the unfalsifiability is coming from strict dependency being true. E.g., a widely-held (but trivial) theory of consciousness is a kind of behaviorism, such that if an entity acts conscious, then it is conscious. But such a theory cannot be falsified, since its predictions could never be any different than experimental inferences (as they are simply a direct restatement of them) and so such a theory contains no scientific information, and so it is trivial. This applies to more than just behaviorism about consciousness. While the term “functionalism” has many different meanings, for the same reason that behaviorist theories of consciousness are trivial, so is any theory of consciousness based solely on input/output. Arguably, this does significantly affect plenty of existent theories, which turn out to be simply behaviorist theories with multiple steps.
@@
“… theorists must resist the temptation to see global accessibility as the cause of consciousness (as if consciousness were some other, further condition); rather, it is consciousness” [dennett2001we].
If Dennett’s ideas were true, this would likely make GNWT a trivial theory of consciousness, since it would be unfalsifiable, because the predictions and inferences strictly depend on the same source. Consciousness just is the information that is globally accessible for report and behavior, and reports and behaviors also just are different expressions of that same information. So there is no situation wherein they could meaningfully diverge. And therefore, the theory is unfalsifiable and trivial (one could see how it gives us minimal scientific information even if it were true).
Often, theories based in the “theater” metaphor have a tendency to fall into strict dependency (sometimes “global workspaces” are described in this way too [baars1997theater]), and too are theories based on identifying only why humans might claim to be conscious (rather than why they are conscious) at high risk of triviality (e.g., theories like the Attention Schema Theory of consciousness [graziano2020consciousness], or any other theory focused on the “meta-problem of consciousness” as an explanandum [chalmers2018meta]).
A significant amount of confusion about the Kleiner-Hoel dilemma was simply not understanding that the dilemma has this second horn of strict dependency. E.g., in one of the only cases of attempting to confront the problem head-on, the criticism ended up being that the substitution argument could not possibly be true, because we did not factor in needing to change the predictions for functionalist theories of consciousness after substitutions [ganesh2020cwars]. But in arguing that predictions must change in the same way as inferences following a substitution, the paper proposes a trivial theory of consciousness (a kind of global input/output functionalism) that falls on the other horn of the Kleiner-Hoel dilemma, thus proving our point that the space is indeed hard to navigate.
In what follows, I will use the Kleiner-Hoel formalism for falsification of theories of consciousness to show several results. In the next section, I show that they provide a disproof of LLM consciousness. After that, I present a case of avoiding the dilemma in humans but not LLMs, indicating that the actual conclusion of the Kleiner-Hoel dilemma is that it disproves consciousness in LLMs but not in humans.
@@ 4 The Proximity Argument
A theory of consciousness $T$ is trivial if there is strict dependency between its predictions and inferences. That is, if the prediction function predpred and inference function infinf are identical (or so approximate that there can be minimal scientific information gained for any possible experiment), or (more generally) such that one cannot be varied without the other, then such theories are unfalsifiable.
Here, I assume that such trivial theories must be false (rather than just trivial), since it would mean there is no scientifically informative theory of consciousness. This assumption makes the language and conclusions easier to follow throughout (e.g., instead of writing “a system for which there are no non-trivial theories of consciousness available” I can say “a non-conscious system”); however, all the conclusions would still hold but with this sort of replacement language if this assumption (which is just that a scientifically-informative theory of consciousness actually exists) were not true.
@@
Following Definition 4.1 of Trivial Theories, a system $s$ is a non-conscious system if, for all theories of consciousness $T$, when $T$ predicts $s$ is conscious, then $T$ must necessarily be trivial.
A lookup table offers an example of a non-conscious system under Definition 4.2. A lookup table $L$ simply implements some input/output function $f:I\rightarrow O$. It has no complexity, no internal dynamics, and is entirely static. Imagine a theory $T$ that claims $L$ has non-trivial consciousness. What could $T$ base its prediction of consciousness on? Certainly, not just the fact that some set of IF-THEN statements exist. The only relevant property for a theory of consciousness when it comes to $L$ is the input or output of the function $f$ itself. However, that’s the identical function that determines all the experimental inference data! Therefore, any $T$ predicting $L$ is conscious must base this prediction solely on $f$, which is identical to the basis of inference, creating strict dependency. Thus any given $T$ is trivial. Thus, $L$ is necessarily a non-conscious system (as long as non-trivial theories of consciousness exist).
@@
For a pair of systems, $s,s^{\prime}\in S$, that share identical input/output (or approximately equivalent input/output, which could be verbal, behavioral, classifications, text, etc.), the substitution distance $d_{S}(s,s^{\prime})$ is determined by the set of properties $\mathcal{P}$ that differ between $s$ and $s^{\prime}$.
As an example of Definition 4.3, consider the well-detailed case of unfolding of a recurrent artificial neural network (RNN) into a feedforward artificial neural network (FNN) that implements the same input/output function, as in [doerig2019unfolding]. In such an unfolding, many properties are lost. E.g., a single-hidden-layer FNN can also approximate any continuous function to arbitrary precision [cybenko1989approximation, hornik1989multilayer]. Yet, it lacks recurrence, temporal dynamics, higher-order structure, broadcasting, depth, etc. On the other hand, some things remain unchanged, like its substrate and activation functions (i.e., it is still an artificial neural network with individual artificial neurons and connections, running on a computer, etc.). The number and scope of these lost properties defines the universal substitution distance. An alternative comparison might be the human brain. Can a human brain be substituted with a functionally-equivalent single-hidden-layer FNN, even one impractically astronomically wide? It is perhaps possible (depending on the notion of possibility)—let us not take sides right now—but the point of this definition is that, if such a substitution were possible, the number of properties “lost” would be much larger (e.g., the brain is biological, contains things like local field potentials, dendritic mechanics, molecular properties, etc.).
@@
For a system $s$ with substitution distance $d_{S}(s,s^{\prime})$ to a non-conscious system $s^{\prime}$, let $\mathcal{P}$ be the set of properties lost under this substitution. A theory of consciousness $T$ is non-trivial only if $T$’s predictions are sensitive to (or grounded in) some subset of $\mathcal{P}$.
This follows naturally from the previous definitions: given a pair of systems, and one is a non-conscious universal substitution, then for a theory of consciousness to be non-trivial it must somehow differentiate the two. If it is a non-trivial theory, it cannot say that the non-conscious system is conscious, so therefore the difference must lie in the properties lost following substitution.
@@
The larger the substitution distance $d_{S}(s,s^{\prime})$ between a system $s$ and a non-conscious system $s^{\prime}$, the larger the space of properties $\mathcal{P}$ that could (at least potentially) ground non-trivial predictions. Conversely, if $d_{S}(s,s^{\prime})$ is small, and so if $s$ and $s^{\prime}$ differ in only a few properties, then there is less room for any of those properties to ground non-trivial predictions. If none of the differing properties can ground non-trivial predictions, then $s$ is also a non-conscious system.
According to the Proximity Principle, given two different systems, one with a very great “substitution distance” from a non-conscious system, and the other with a small “substitution distance” from a non-conscious system, then in the latter case there are much fewer properties that could make a difference when it comes to claims of consciousness.
@@
Any non-trivial theory of consciousness that predicts contemporary LLMs are conscious must base its predictions entirely on properties constrained within the substitution distance between a lookup table and a single-hidden-layer FNN, combined with the substitution distance between a single-hidden-layer FNN and an LLM.
@@
Step 1. Let $M$ be a static LLM at inference time, let $N$ be a similarly static single-hidden-layer feedforward neural network implementing the same input/output function, and let $L$ be a lookup table implementing the same input/output function.
Step 2. By the universal approximation theorem, any function computed by $M$ can be precisely approximated by a sufficiently wide static single-hidden-layer network $N$. By definition, any function computed by $N$ can in turn be represented as a lookup table $L$ (for finite/discretized inputs, which LLMs have, given fixed vocabulary and context length).
Step 3. Therefore we have a chain of universal substitutions preserving input/output:
@@
Step 4. By Definition 4.2, $L$ is a non-conscious system. By the Proximity Principle (Proposition 4.5), any non-trivial theory $T$ predicting $M$ is conscious must base its predictions solely on properties in $d_{S}(M,L)\subseteq d_{S}(M,N)\cup d_{S}(N,L)$ . ∎
@@
If no property within this constrained space can ground non-trivial predictions for a theory of consciousness, then LLMs are non-conscious systems.
At this point in the chain of reasoning, we should already have strong reasons to question LLM consciousness, due to how significant this constraint is. Certainly, there are some details like scaffolding, interfaces, what type of contemporary LLM, etc., that I am leaving aside. However, at least when it comes to the core “bare bones” of a baseline contemporary LLM (like, say, the original GPT-4, with no router or other complications), any claim to consciousness is confined to its distance to a static single-hidden-layer approximation of its input/output (via the universal approximation theorem [hornik1989multilayer]). And note: Here, the substitution of a single-hidden-layer FNN is another kind of artificial neural network that computes things in much the same way, in that it has similar activation functions, is also instantiated on a computer, has identical input/output, and so on.
That is, if we just stopped here at Corollary 4.7, LLM consciousness would, at minimum, be highly constrained. It must require something lost in that chain of universal substitutions (like, say, extra layers). If one still holds that a single-hidden-layer static FNN network could be non-trivially conscious, then there is also the distance to the provably-trivial lookup table to consider (there, the only relevant property appears to be compression, which is also not used for most theories of consciousness—following this logic, see Corollary 4.10 for a proof that single-hidden-layer static FNNs are non-conscious). If a theory of consciousness cannot specify what is important and lost in this substitution chain, then it cannot possibly claim that LLMs are conscious.
However, one can go further, because this so far has only discussed the second horn of the Kleiner-Hoel dilemma: the relationship between triviality and strict dependency. If we add back in the first horn to this transitive comparison, the proof rules out baseline LLM consciousness entirely.
@@
Step 1. The substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to preserve input/output. Since inferences for entities like LLMs are derived from input/output functions (not from, e.g., how many layers they have), all inference data is fixed across the chain.
Step 2. Let $\rho$ be any property in $d_{S}(M,L)\subseteq d_{S}(M,N)\cup d_{S}(N,L)$. By definition of substitution distance, $\rho$ varies across the chain.
Step 3. For any theory $T$ grounding predictions on $\rho$: since $\rho$ varies while inference data is by definition fixed, predictions are necessarily independent from inferences with respect to $\rho$.
Step 4. By the Substitution Argument, $T$ is a priori falsified.
Step 5. The only alternative is to ground predictions in the input/output structure itself, which creates strict dependency with inferences, rendering $T$ trivial.
Step 6. Therefore, for any given $T$ it either: (a) grants consciousness to contemporary baseline LLMs, but does not for trivially non-conscious systems, but then $T$ is a priori falsified by mismatches in prediction when non-conscious systems are substituted in; or (b) grants consciousness to LLMs and trivially non-conscious systems, but then the theory must apply in the latter case, and so must be strictly dependent and trivial, being based only on input/output. ∎
An alternative summary of Proposition 4.8 is to say that the substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to fix input/output, which essentially exhausts all relevant inference data (especially in light of Corollary 4.10, which proves static single-hidden-layer FNNs are non-conscious), since there would have to be a priori neutral and non-empirical and non-theory-based reasons to assume LLM consciousness, which somehow do not apply to single-hidden-layer FNN consciousness, and to such a degree that it (somehow) overrides the pathological mismatches with a non-conscious system. Therefore, any property that varies across the chain should be independent of inferences. This eliminates the possibility of lenient dependency within this specific substitution space.
@@
An example of how the Proximity Argument works can be seen by zooming into one link in the chain from lookup tables to single-hidden-layer FNNs. Specifically, we can make the same form of argument there, using a specific property within the substitution distance (and the only one available that can substantially ground a theory of consciousness).
Corollary 4.10 (Single-hidden-layer FNNs are non-conscious systems).
Examining a single step in the Proximity Principle: the substitution distance $d_{S}(N,L)$ between a lookup table $L$ and a single-hidden-layer FNN $N$ implementing the same input/output function can be entirely summarized as compression, and compression can be varied via universal substitutions, and therefore single-hidden-layer FNNs are non-conscious systems.
@@
Step 1. A lookup table over $n$ discrete inputs can be implemented as a maximally wide single-hidden-layer FNN (one hidden unit per possible input/output).
@@
Step 2. Substitute in a narrower network, but with identical input/output. Given that input/output is fixed, the only difference is compression of the original function.
@@
Step 3. Any theory $T$ that grounds predictions in compression can therefore yield independent predictions across these networks, all while inferences (derived from input/output) remain fixed—thus creating the pathological mismatches that falsify $T$.
@@
Step 4. Therefore, only the original input/output remains to ground $T$, and so by Definition 4.2 of non-conscious systems, single-hidden-layer FNNs are non-conscious systems. ∎

### v2

@@ 3 A Formal Framework for Falsifying Theories of Consciousness
What universal substitutions provide is a means of a priori falsification. Essentially, substituting them into an experimental setup for the original would radically alter or complicate the predictions of a theory of consciousness, while the outward and immediate inferences (such as report or output) would not change at all (other than irrelevant considerations for testing a theory of consciousness, like say, exact timing of reports).
There are two main classes of universal substitutions: causal-structure substitutions and algorithmic substitutions.
  • Causal-structure substitutions accomplish a particular input/output function via a different system with different internal workings. This includes lookup tables that implement the same input/output function of a system directly [63], or static FNNs that implement the same input/output (or approximate it arbitrarily well) [31]. This might include potentially more exotic options, like substitutions based on the AIXI model of universal intelligence [40] or even cases of a brain in quantum superposition (a “Schrödinger’s Zombie”) [18].
  • Functional substitutions accomplish any particular input/output function via a different algorithm or program with different workings. For instance, there are various ways to sort (bubble sort, merge, etc.) that can all transition an unsorted list to a sorted one.
It had been hoped that issues of falsification like the Unfolding Argument would only ever impact “causal structure” theories like IIT [31]. Alternatively, some have speculated that if this form of argument impacts functionalist theories, it would form a reductio and actually argue against the Kleiner-Hoel dilemma being a problem [72]. However, that its widespread negative result is too strong would only be true if the Kleiner-Hoel dilemma provably cannot be navigated by any theory whatsoever. If a theory (or a class of theories) does navigate it, we cannot say that the requirement is too strong: rather, it is precisely strong enough to substantially constrain consciousness research and point it in the right direction.
For these purposes, in what follows I show precisely how the Kleiner-Hoel dilemma impacts all functionalist theories of consciousness too. Instead of substituting systems, we can substitute programs. Specifically, the shortest program.
Theorem 3.1 (Kolmogorov Substitution Theorem).
Following [52], for any computable function $f$ there exists a shortest program $K(f)$ implementing $f$ (even if it is not known how to identify it). This $K(f)$ can always be used as a universal substitution for any other program $N$ implementing $f$, and so will cause mismatches and falsify a given theory of consciousness that targets the program $N$ as the basis of its predictions. Alternatively, if a theory of consciousness has the same predictions for all programs implementing $f$, including $K(f)$, then the predictions are entirely determined by $f$ itself, not the original program implementing it, since $K(f)$ is solely a function of $f$. In that case, predictions and inferences have the same source and cannot vary, leading to triviality.
Proof.
Step 1. For any computable $f$, $K(f)$, the shortest program implementing $f$, exists, and is itself solely a function of $f$.
Step 2. Let $N$ be any program computing $f$ that is the target of the predictions of some theory of consciousness. $N$ and $K(f)$ share input/output (defined by $f$).
Step 3. Any theory that assigns different predictions about consciousness to $N$ (the actual program under question) and $K(f)$ (its universal substitution) is falsified, since this implies empirical mismatches.
Step 4. Any theory that always shares predictions between $N$ and $K(f)$ is basing those predictions on properties of $K(f)$ alone. Since predictions based on $K(f)$ can be derived entirely without any knowledge of program $N$, this renders program $N$ irrelevant to the given theory of consciousness. Therefore, any such theory collapses into triviality. ∎
Many theories could be shown to have problems with falsification in this light. Consider just one: Higher Order Thought (HOT) theory, a popular theory of consciousness [19]. Any exact ranking or rating of the different representations should differ in its predictions about $N$ vs. $K(f)$. But if somehow not, then $N$ is irrelevant and only the input/output and its subsequently determined $K(f)$ is needed for the theory’s predictions, rendering it trivial. The same goes for almost all theories of consciousness that are more focused on functionalist aspects, like “representation” or “computation.”
The overall structure of this proof also demonstrates how the Kleiner-Hoel dilemma has two horns. Here, if substituting in the shortest program that computes the function $f$ makes a theory’s predictions change, then it creates mismatches: A kind of “a priori falsification.” However, on the other hand, if the predictions of a theory of consciousness do not change, then they must be strictly dependent on the experimental inferences observed during testing a theory of consciousness, rendering that theory trivial (here, this can be clearly seen by any details of the original program $N$ no longer mattering at all if a theory’s predictions are always the same for $K(f)$ too). This would be a case of “strict dependency” wherein a predictions can be entirely derived by the inferences (in the terminology of Theorem 3.1, inferences are entirely determined by $f$).
@@
Here and throughout, such strictly-dependent-on-inferences theories of consciousness are referred to as “trivial” to clearly distinguish that the source of the unfalsifiability is coming from strict dependency holding true. E.g., a widely-held (but trivial) theory of consciousness is a kind of behaviorism, such that if an entity acts conscious, then it is conscious. But such a theory cannot be falsified, since its predictions could never be any different than experimental inferences (as they are simply a direct restatement of them) and so such a theory contains no scientific information, and so it is trivial. This applies to more than just behaviorism about consciousness. While the term “functionalism” has many different meanings, for the same reason that behaviorist theories of consciousness are trivial, so is any theory of consciousness based solely on input/output. Arguably, this does significantly affect plenty of existing theories, which turn out to be simply behaviorist theories with multiple steps.
@@
“… theorists must resist the temptation to see global accessibility as the cause of consciousness (as if consciousness were some other, further condition); rather, it is consciousness” [29].
If Dennett’s ideas were true, this would make GNWT a trivial theory of consciousness, since it would be unfalsifiable, because the predictions and inferences strictly depend on the same source. Consciousness just is the information that is globally accessible for report and behavior, and reports and behaviors also just are different expressions of that same information. So there is no situation wherein they could meaningfully diverge. And therefore, the theory is unfalsifiable and trivial (one could see how it gives us minimal scientific information even if it were true—trivial theories are fundamentally unsatisfying).
Often, theories based in the “theater” metaphor have a tendency to fall into strict dependency (sometimes “global workspaces” are described in this way too [10]). Theories based on identifying only why humans might claim to be conscious (rather than why they are conscious) are at high risk of triviality (e.g., theories like the Attention Schema Theory of consciousness [35], or any other theory focused on the “meta-problem of consciousness” as an explanandum [23]).
A significant amount of confusion about the Kleiner-Hoel dilemma was simply not understanding that the dilemma has this second horn of strict dependency. E.g., in one of the only cases in the literature of attempting to confront the problem head-on, the criticism ended up being that the substitution argument could not possibly be true, because we did not factor in needing to change the predictions for functionalist theories of consciousness after substitutions [34]. But in arguing that predictions must change just the same way as inferences following a substitution, the paper proposes a trivial theory of consciousness (a kind of global input/output functionalism) that falls on the other horn of the Kleiner-Hoel dilemma, thus proving our point that the space is indeed hard to navigate.
In what follows, the Kleiner-Hoel formalism for falsification of theories of consciousness is used to show several results. In the next section, I show that it provides a disproof of LLM consciousness. After that, I present a case of avoiding the dilemma in humans but not LLMs, indicating that the actual conclusion of the Kleiner-Hoel dilemma is that it disproves consciousness in LLMs but not in humans.
@@ 4 The Proximity Argument
A theory of consciousness, $T$, is trivial if there is strict dependency between its predictions and inferences. That is, if the prediction function, predpred, and inference function, infinf, are identical (or if the two are so approximate there can only be minimal scientific information gained for any possible experiment). An example of triviality is if infinf cannot be varied without predpred also being varied, or both stem from the same source of data, then such a theory is unfalsifiable since mismatches are impossible.
Here, I assume that trivial theories of consciousness must be false (rather than just trivial), since otherwise there is no scientifically informative theory of consciousness. This assumption makes the language and conclusions easier to follow throughout (e.g., instead of writing “a system for which there are no non-trivial theories of consciousness available” I can say “a non-conscious system”); however, all the conclusions would still hold but with this sort of replacement language if this assumption (which is just that a scientifically-informative theory of consciousness actually exists) were not true.
@@
Following Definition 4.1 of Trivial Theories: A system $s$ is a non-conscious system if, for any theory of consciousness $T$, whenever $T$ predicts $s$ is conscious, $T$ must necessarily be trivial.
A lookup table offers an archetypal example of a non-conscious system under Definition 4.2. Any lookup table $L$ simply implements a input/output function $f:I\rightarrow O$. Imagine a theory $T$ that predicts $L$ has non-trivial consciousness. What could $T$ base its prediction of consciousness on? $L$ is not internally complex, lacks all dynamics, is entirely static, has no memory capabilities, and there is no information flow across any parts or modules. Certainly, a theory cannot make meaningful predictions of consciousness based on merely the fact that some set of IF-THEN statements exist. The sole relevant property for a theory of consciousness pertaining to $L$ is the input/output of the function $f$ itself. However, that’s the same function that determines all the experimental inference data! Therefore, any $T$ predicting $L$ is conscious must base this prediction solely on $f$, which is identical to the basis of inference, creating strict dependency. Thus any given $T$ is trivial. Thus, $L$ is necessarily a non-conscious system (as long as, per definition 4.1, non-trivial theories of consciousness exist).
@@
For a pair of systems, $s$ and $s^{\prime}$, that share identical input/output (or approximately equivalent input/output, wherein this could describe behavior or responding to text, etc.), the substitution distance $d_{S}(s,s^{\prime})$ is determined by the space of properties $\mathcal{P}$ that differ between $s$ and $s^{\prime}$.
As an example of Definition 4.3, consider the well-detailed case of unfolding a recurrent artificial neural network (RNN) into a feedforward artificial neural network (FNN) that implements the same input/output function, as in [31]. In such an unfolding, many properties are lost. E.g., while a single-hidden-layer FNN can approximate any continuous function to arbitrary precision [28, 39], it lacks recurrence, temporal dynamics, higher-order structure, broadcasting, depth, etc. On the other hand, some things remain unchanged, like its substrate and activation functions (i.e., it is still an artificial neural network with individual artificial neurons and connections, running on a computer, etc.). The scope of these lost properties defines the substitution distance (Definition 4.3).
An alternative comparison might be the human brain. Can a human brain be substituted with a functionally-equivalent single-hidden-layer FNN, even one impractically astronomically wide? It is perhaps possible (depending on the notion of possibility)—let us not take sides right now—but the point of this definition is that, if such a substitution were possible, the number of properties “lost” would be much larger (e.g., the brain is biological, contains things like local field potentials, dendritic mechanics, molecular properties, etc.). And the same goes for substituting a human brain with the shortest program capable of carrying out its input/output function (as in Theorem 3.1); much more would be lost than just substituting one computer program for another.
@@
For a system $s$ with substitution distance $d_{S}(s,s^{\prime})$ to a non-conscious system $s^{\prime}$, let $\mathcal{P}$ be the set of properties lost following this substitution. A theory of consciousness, $T$, is non-trivial only if $T$’s predictions are sensitive to (or grounded in) some subset of $\mathcal{P}$, since by Definition 4.2, for non-conscious systems only trivial theories are possible.
This follows naturally from the previous definitions: given a pair of systems, and one is a non-conscious substitution, then for a theory of consciousness to be non-trivial it must somehow differentiate the two. If it is a non-trivial theory of consciousness, it cannot predict that the non-conscious system is conscious, so therefore the difference must lie in the properties lost compared to the non-conscious substitution.
@@
The larger the substitution distance $d_{S}(s,s^{\prime})$ between a system $s$ and a non-conscious system $s^{\prime}$, the larger the space of properties $\mathcal{P}$ that could (at least potentially) ground non-trivial predictions. Conversely, if $d_{S}(s,s^{\prime})$ is small, and so if $s$ and $s^{\prime}$ differ in only a few properties, then there is a small space of properties that could ground non-trivial predictions. If none of the differing properties can ground non-trivial predictions, then $s$ is also a non-conscious system.
According to the Proximity Principle, given two different systems, one with a large “substitution distance” from a non-conscious system, and the other with a small “substitution distance” from a non-conscious system, then in the latter case there are much fewer properties that could make a difference when it comes to predictions of consciousness.
@@
Any non-trivial theory of consciousness that predicts contemporary LLMs are conscious must base its predictions entirely on properties constrained to the substitution distance between a lookup table and a static single-hidden-layer FNN, combined with the substitution distance between a static single-hidden-layer FNN and an LLM.
@@
Step 1. Let $M$ be a deployed static LLM, let $N$ be a static single-hidden-layer feedforward neural network implementing the same input/output function as $M$, and let $L$ be a lookup table also implementing the same input/output function.
Step 2. By the universal approximation theorem, any input/output function (like text responses) computed by $M$ can be approximated arbitrarily well by a sufficiently-wide static single-hidden-layer network $N$. By definition, any finite function computed by $N$ can in turn also be implemented as a lookup table, $L$.
Step 3. Therefore there exists a chain of universal substitutions preserving input/output:
@@
Step 4. By Definition 4.2, $L$ is a non-conscious system. By the Proximity Principle (Proposition 4.5), any non-trivial theory $T$ predicting $M$ is conscious must base its predictions solely on properties within the space defined by $d_{S}(M,L)\subseteq d_{S}(M,N)\cup d_{S}(N,L)$. ∎
@@
If no property within this space constrained by this universal substitution chain can ground non-trivial predictions for a theory of consciousness, then LLMs are non-conscious systems.
At this point in our reasoning, we should already have strong reasons to question LLM consciousness, simply due to how significant this constraint is. Certainly, there are some details like scaffolding, interfaces, if it is deployed, and what type of contemporary LLM, etc., that I am leaving aside. However, at least when it comes to the core “bare bones” of a baseline contemporary LLM (like, say, the original GPT-4, with no router or other complications), any claim to consciousness is confined to its distance to a static single-hidden-layer approximation of its input/output (which is guaranteed by the universal approximation theorem [39]). And note: Here, the substitution of a static single-hidden-layer FNN is another kind of artificial neural network that computes input/output in much the same way, in that it has similar activation functions, is also instantiated on a computer, has identical input/output, and so on. So if we just stopped here at Corollary 4.7, LLM consciousness would, at minimum, be highly constrained. It must require something lost in that chain of universal substitutions (like, say, extra layers).
If one still holds that a single-hidden-layer static FNN network could be non-trivially conscious, then there is also the distance to the provably-trivial lookup table to consider (there, the only relevant property appears to be compression, which is also not a good target for theories of consciousness—following this logic, see Corollary 4.10 for a proof that static single-hidden-layer static FNNs are non-conscious). If a theory of consciousness cannot specify what is important and lost along this substitution chain, then it cannot claim that LLMs are conscious.
However, one can go further, because this so far has only discussed the second horn of the Kleiner-Hoel dilemma: the relationship between triviality and strict dependency. If we add back in the first horn of mismatches following substitutions to this transitive chain, the proof rules out baseline LLM consciousness entirely.
@@
Step 1. The substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to preserve input/output. Since empirical inferences about consciousness for entities like LLMs are derived from input/output functions (not from, e.g., how many layers they possess), all inference data is fixed across the chain.
Step 2. Let $\rho$ be some specific property in $d_{S}(M,L)\subseteq d_{S}(M,N)\cup d_{S}(N,L)$. By Definition 4.3 of Substitution Distance, assume that $\rho$ varies across the chain.
Step 3. For any theory $T$ that bases its predictions on $\rho$: since $\rho$ varies while inference data is by definition fixed, predictions are necessarily independent from inferences with respect to $\rho$. This necessarily leads to mismatches, meaning that $T$ is a priori falsified.
Step 4. The only alternative is if $\rho$ does not vary, in which case, $T$ must necessarily ground predictions solely in the input/output structure itself. This creates a situation of strict dependency, rendering $T$ trivial.
Step 5. Therefore, for any given theory of consciousness $T$ it either: (a) grants consciousness to contemporary baseline LLMs, but does not for trivially non-conscious systems, yet then $T$ is a priori falsified by mismatches in prediction when non-conscious systems are substituted in for LLMs; or (b) grants consciousness to LLMs and non-conscious systems, but then the theory must apply in the latter case, and so must be trivial. ∎
An alternative summary of Proposition 4.8 is to say that the substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to fix input/output, which essentially exhausts all relevant inference data (especially in light of Corollary 4.10, which proves static single-hidden-layer FNNs are non-conscious), since there would have to be a priori neutral and non-empirical and non-theory-based reasons to assume LLM consciousness, which somehow do not apply to single-hidden-layer FNN consciousness, and to such a degree that it (somehow) overrides the pathological mismatches with a non-conscious system. Therefore, any property that varies across the chain should be independent of inferences. This eliminates the possibility of overcoming the Kleiner-Hoel dilemma within this substitution chain.
@@
An example of how the Proximity Argument works can be seen by zooming into one link in the chain from lookup tables to single-hidden-layer FNNs. Specifically, we can make the same form of argument there, using a specific property, compression, that is within the substitution distance (and the only one available substantial enough to ground a theory of consciousness).
Corollary 4.10 (Static Single-hidden-layer FNNs are non-conscious systems).
Examining a single step in the Proximity Principle: consider the substitution distance $d_{S}(N,L)$ between a lookup table, $L$, and a static single-hidden-layer FNN, $N$, implementing the same input/output function. This distance can be entirely summarized as the degree of compression of the input/output function, and since compression can be varied via universal substitutions, therefore single-hidden-layer FNNs are non-conscious systems.
@@
Step 1. A lookup table with $n$ discrete inputs can be implemented as a maximally wide static single-hidden-layer FNN (with one hidden unit per possible input/output).
@@
Step 2. Substitute in a narrower network with identical input/output. Since that input/output is fixed, the only difference is in the degree of compression in the implementation.
@@
Step 3. Any theory $T$ that grounds its predictions of consciousness in compression can therefore yield independent predictions across networks, all while inferences (necessarily derived from input/output) remain fixed—thus creating the pathological mismatches that falsify $T$.
@@
Step 4. Therefore, only the input/output function itself remains to actually ground $T$, and so by Definition 4.2 of non-conscious systems, static single-hidden-layer FNNs are non-conscious systems. ∎
Regarding Corollary 4.10, it is notable that $K(f)$ could be used instead, representing the point of maximal compression (while lookup tables represent the point of minimal compression).